How much applied force is needed to given an 8-N object (a) an upward acceleration 
of 2 m/s2 and (b) a downward acceleration of 2 m/s2
? (c) In what direction must the latter 
force act?

Grade:12th pass

1 Answers

CHIRAG HIWAL
4 years ago
(a) here we have force of 8N in downward direction and acceleration of 2m/s^2 in uppward. let we apply force F in upward direction. 
 Here,
mg=8N
mass of obj is 8/9.8 =0.816kg
So, by newton’s law of motion 
ma=F-mg
i.e. F= ma+mg
F=0.816*2+8=9.63N
 
(b) acc. 2m/s^2 uppward 
similarly,
ma=mg-F
i.e. F=8-2*0.816=6.37N
 
© when we apply both forces the block move upward.